OpenStudy (anonymous):

solve the following on the interval (0, 306) 2sec^2x-secx+3=0

OpenStudy (anonymous):

what do you mean by that?

OpenStudy (anonymous):

No, i started out by simplifying it but i don't think i did it right

OpenStudy (anonymous):

I did this: 2sec^2-secx=3 divided 2 to the other side took the square root of sec^2 and 3/2

OpenStudy (anonymous):

now I'm a little stumped

OpenStudy (anonymous):

2sec^2x-secx=t

OpenStudy (anonymous):

or t+3=0

OpenStudy (loser66):

oops, I am sorry, you have a complex number

OpenStudy (anonymous):

Did i do it wrong in the first place ?

OpenStudy (loser66):

please recheck your problem, first off , the interval (0, 306) ??? or (0, 360) ?? second the equation 2sec^2 (x) - sec(x) +3 =0 or sec ^2 (x) - 2sec (x) +3 =0??

OpenStudy (anonymous):

(0,360) oops.

OpenStudy (loser66):

and the equation?

OpenStudy (anonymous):

2sec^2x-secx+3=0

OpenStudy (anonymous):

I'm supposer to get multiple answers for x if that helps

OpenStudy (loser66):

wow, I am stuck at complex number. http://www.wolframalpha.com/input/?i=2sec^2x+-sec+x+%2B3+%3D0

OpenStudy (anonymous):

oh no that doesn't look so good, it's ok thank you for trying i really appreciate it!

OpenStudy (loser66):

hehehe, I tried and stuck . Then, I looked for the reason, it gave me the reason, hehehe... I failed!! hehehehe I am sorry for being helpless

OpenStudy (anonymous):

You weren't helpless! at least you tried ! Thank you- i'll ask my teacher tomorrow howl to solve it

OpenStudy (anonymous):

you need to sub tract the 3 then factor to get so you can het something that looks like this (a+1)(a-1) you get what im saying ?